find the measure of angle BAC if the measure of angle BAD=68, the measure of angle BAC=2x-3 and the measure of angle CAD=4x+5

Answer by Theo(2967)

looks like angle BAC plus angle CAD = angle BAD.
angle BAD = 68
angle BAC = 2x-3
angle CAD = 4x+5
we have:
2x-3 + (4x+5) = 68
which becomes:
2x - 3 + 4x + 5 = 68
which becomes
6x + 2 = 68
subtract 2 from each side of the equation to get:
6x = 66
divide both sides of the equation by 6 to get:
x = 11
-----
if x = 11, then:
angle BAC = 2x - 3 = 22 - 3 = 19
angle CAD = 4x + 5 = 44 + 5 = 49
19 + 49 = 68
since this is true, the answer is good.
answer to your question is:
measure of angle BAC = 19 degrees.